OpenStudy (potatoe):

set up the partial fraction: integral of [dx/x(x^2-1)]

OpenStudy (lgbasallote):

\[\LARGE \frac{1}{x(x^2 - 1)} \implies \frac{1}{x(x+1)(x-1)}\] does that help?

OpenStudy (potatoe):

ok so a/x..+..b/(x+1)..+..c/(x-1)?

OpenStudy (lgbasallote):

you got it

OpenStudy (potatoe):

thanks

OpenStudy (lgbasallote):

\[\frac{A}{x} + \frac{B}{x+1} + \frac{C}{x-1}\] welcome ^_^